Marvel Parts, Inc., manufactures auto accessories. One of the companys products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 2.850 hours each month to produce 1,900 sets of covers. The standard costs associated with this level of production are: Per Set of Total Covers Direct materials Direct labor Variable manufacturing overhead $42,560 $22.40 $51,300 27.00 (based on direct labor-hours) $ 6,840 3.60 $53.00 During August, the factory worked only 2,800 direct labor-hours and produced 2,000 sets of covers. The following actual costs were recorded during the month: Per Set Total of Covers Direct materials (12,000 yards) $45,600 $22.80 $49,e00 24.50 Variable manufacturing overhead $ 7,e00 3.50 $50.80 Direct labor At standard, each set of covers should require 5.6 yards of material. All of the materials purchased during the month were used in productionAt standard, each set of covers should require 5.6 yards of material. All of the materials purchased during the month were used in production Required . Compute the materials price and quantity variances for August. 2. Compute the labor rate and efficiency variances for August. . Compute the variable overhead rate and efficiency variances for August. (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance). Input all amounts as positive values.) Materials price variance Materials quantity variance 2 Labor rate variance Labor efficiency variance 3 Variable overhead rate variance Variable overhead efficiency variance
Answer #1

1) Material price variance = (4*12000-45600) = $2400 F

Material quantity variance = (2000*5.6-12000)*4 = 3200 U

2) Labor rate variance = (18*2800-49000) = 1400 F

Labor efficiency variance = (2000*1.5-2800)*18 = 3600 F

3) Variable overhead rate variance = (2.4*2800-7000) = 280 U

Variable overhead efficiency variance = (3000-2800)*2.4 = 480 F